ROTATIONAL CONSTANTS FOR THE HIGH ENERGY ROTAMER OF 1,3-BUTADIENE

\underline{Yu. N. PANCHENKO}, Department of Chemistry, Moscow State University, Moscow 119899, Russia.

A region of the IR spectrum of 1,3-butadiene containing only a band of the s-gauche rotamer was predicted by quantum mechanical calculations. Investigation of the IR spectrum using Fourier transform spectroscopy in the gas phase revealed the presence of a band at 749.22 (20) cm^{−1} possessing the typical b-type contour consistent with the predicted A species, (C_{2} symmetry) for the ν_{12} band (=CH_{2} twist) of the s-gauche rotamer. Rotational analysis of this band on the spectrum recorded for the first time at a resolution of 0.64 cm^{−1} yielded (A′ − B̃′) = 0.4455 (25) cm^{−1} and (A′′ − B̃′′) = 0.4478 (27) cm^{−1} ^{a}. The extremely complicated character of the high resolution spectrum of this band suggests also non-planar s-gauche structure.

aG. R. De Maré, Yu. N. Panchenko and J. Vander Auwera, to be published.